# Unstable entropy for Anosov diffeomorphisms on the $3$-torus

### v1

@@ 3. Construction of a Margulis system of measures
which shows that $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ has full support in each leaf and is finite on compact sets. The system $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ is measurable. We use that $\mu^{u}\in\mathcal{X}$, which means that there is a sequence
$\mu_{m}:=\sum_{i=1}^{n_{m}}\underbrace{t_{i,m}}_{>0}\nu^{n_{m}}\in\mathrm{Conv}\big{\{}\nu^{n}:=e^{-nh^{u}_{\mathrm{top}}(f)}(f^{*})^{n}\mathrm{Vol}_{\mathcal{W}^{u}}\mid n\in\mathbb{N}\big{\}},\ \lim_{m\to+\infty}\mu_{m}=\mu^{u}.$
Let $B$ be a foliation chart and let $\phi\in C_{c}^{\infty}(B,\mathbb{R})$. For any $x\in B$, one has
$\mu^{u}_{x}(\phi|_{|\mathcal{W}^{u}_{B}(x)})=\lim_{m\to+\infty}(\mu_{m})_{x}(\phi|_{|\mathcal{W}^{u}_{B}(x)}).$
Note that for any $n\geq 0$, the mapping $x\mapsto(\nu^{n})_{x}(\phi|_{|\mathcal{W}^{u}_{B}(x)})$ is continuous (hence measurable). This means that $x\mapsto(\mu_{m})_{x}(\phi|_{|\mathcal{W}^{u}_{B}(x)})$ is also measurable as a convex combination of measurable functions. Since the limit of a sequence of measurable functions is measurable, we deduce that $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ is a measurable system. ∎
@@
Having constructed a measurable system of $u$-measures, we now characterize $u$-MMEs by their conditionals along the strong unstable foliation. We recall the following result due to Buzzi, Fisher and Tazhibi [9], see also [38, Proposition 5.3].
@@ 4.1. Ruelle resonances
In this subsection, we recall some facts about the functional approach for Anosov diffeomorphisms. For a detailed introduction to the subject, we refer to [2]. For $0\leq k\leq 3$, define the vector bundle of smooth $k$-forms $\Omega_{k}=C^{\infty}(\mathbb{T}^{3};\Lambda^{k}T^{*}(\mathbb{T}^{3})).$ The diffeomorphism $f$ acts on $\Omega_{k}$ by pushforward:
@@
For an Anosov diffeomorphism, one can associate to $f_{*,k}$ a discrete spectrum $\mathrm{Res}_{k}(f)$, the Pollicott-Ruelle resonances, by making it act on specially designed anisotropic spaces, see for instance [5, 3, 18, 16].
@@
where $D^{\prime}_{k}(\mathbb{T}^{3})$ denotes the space of distributional currents of degree $k$, i.e., the dual space of $\Omega_{3-k}$, see [29, Chapter 10.1]. We will write $(\cdot,\cdot)_{\Omega_{k} \times\mathcal{D}^{\prime}_{3-k}}$ for the pairing between $\Omega_{k}$ and $\mathcal{D}^{\prime}_{3-k}$. More precisely, by the work of Faure, Roy and Sjöstrand [16], there exists a family of Hilbert spaces $(\mathcal{H}^{s}(\mathbb{T}^{3};\Omega_{k}))_{s>0}$ such that:
  • the space $\Omega_{k}$ is densely included in $\mathcal{H}^{s}(\mathbb{T}^{3};\Omega_{k})$ (see [29, Lemma 9.1.13]) and one has $H^{s}\subset\mathcal{H}^{s}\subset H^{-s}$, where $H^{s}$ is the usual $L^{2}$-Sobolev space of order $s$ (see [29, Lemma 9.1.14]);
@@
Applying the result to $f^{-1}$, we obtain a family of Hilbert spaces $(\mathcal{H}^{-s}(\mathbb{T}^{3},\Omega_{k}))_{s>0}$, dual to $(\mathcal{H}^{s}(\mathbb{T}^{3},\Omega_{k}))_{s>0}$ on which the resolvent of the pullback of $f$ has a meromorphic extension, see [16]. We will write $(\cdot,\cdot)_{\mathcal{H}^{s} \times\mathcal{H}^{-s}}$ for the pairing between $\mathcal{H}^{s}$ and $\mathcal{H}^{-s}$.
@@
where $R_{k}^{H}(\lambda)$ is the holomorphic part near $\lambda_{0}$. The generalized resonant states are the elements in the range of the spectral projector:
@@
where $\mathrm{WF}(u)$ denotes the wavefront set of a current $u$, see [23, Chapter VIII]. The resonant states are defined as
@@ 4.2. Zeta function
The zeta function $\zeta_{f}$ has a meromorphic extension (actually rational extension in our case) to the complex plane. This is a special case of the works of Liverani-Tsujii [30] and Baladi-Tsujii [4]. In fact, for an Anosov diffeomorphism on a torus, one can compute explicitly $\zeta_{f}$ in terms of its action on the homology $A$.
@@ 4.3. Dynamical determinants
We will need the following result, see for instance [2, Theorem 6.2]. For any $k\in\llbracket 0,3\rrbracket$, the dynamical determinant $\mathcal{D}_{f,k}$ admits a holomorphic extension to $\mathbb{C}$. Moreover,
@@
We recognize the Ruelle zeta function with potential $J^{c}_{f}$. Since $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ is conjugate to a linear automorphism of $\mathbb{T}^{3}$, it is topologically mixing. By [35, Theorem 1], there is a $R=R(J^{c}_{f})>e^{-P(J^{c}_{f})}$ such that $\zeta_{f,J^{c}_{f}}$ has no zero in $\overline{B}(0,R)$ except for a simple zero at $e^{-P(J^{c}_{f})}$. This shows that $e^{P(J^{c}_{f})}$ is a resonance of algebraic multiplicity equal to $1$ and that $e^{P(J^{c}_{f})}$ is the only resonance on the critical circle. ∎
@@
see [2, Equation $(6.38)$]. Using (4.9), we also show the following.
@@
We show that the trace of the spectral projector at the first resonance is equal to $\mu_{J^{c}_{f}}$.
@@
Let $g\in C^{\infty}(\mathbb{T}^{3},\mathbb{R}_{+})$ be a smooth and positive function. The Guillemin trace formula (see for instance [4, Proposition 6.3]) gives for any $m\geq 0$,
@@
This means that (4.17) is Bowen’s formula for the equilibrium state (see for instance $\mu_{J^{c}_{f}}$ [21, Theorem 20.3.7]) and this concludes the proof. ∎
@@
Let $\omega\in\mathcal{H}^{s}$, then there is $c_{\omega}\in\mathbb{R}$ such that $\Pi_{0}(\omega)=c_{\omega}\theta$. We obtain (4.18) by applying the previous convergence to $\omega$. The proof of (4.19) is analogous. ∎
@@
By [21, Theorem 20.3.9], $\mu_{J^{c}_{f}}$ is a measure of maximal entropy if and only if $J^{c}_{f}$ is cohomologous to a constant. By the Livsic theorem [21, Theorem 19.2.1], this holds if and only if for any periodic point $p$ of period $n(p)$, one has
@@
The fact that the constant has to be equal to $\lambda_{c}$ follows from [17, Corollary 2.6]. This holds if and only if $E_{u}$ and $E_{s}$ are jointly integrable by [17, Theorem 5.1]. ∎
@@
We note that by Corollary 1.2 and Corollary 4.2, the system of $u$-measures $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ coincides (up to a constant rescaling) with the conditionals $(\mu^{u}_{J^{c}_{f},x})_{x\in\mathbb{T}^{3}}$ of $\mu_{J^{c}_{f}}$ along $\mathcal{W}^{u}$. We first start by expressing the $cs$-conditionals $(\mu^{cs}_{J^{c}_{f},x})_{x\in\mathbb{T}^{3}}$ using the resonant state $\theta$. By (1.6), one has $\mathrm{WF}(\theta)\subset E_{cu}^{*}$ and for any $x\in\mathbb{T}^{3}$, $\mathrm{WF}([\mathcal{W}^{cs}(x)])\subset E_{cs}^{*}$, where $\mathcal{W}^{cs}(x)]$ denotes the integration current on $\mathcal{W}^{cs}(x)$, see for instance [29, Example 4.1.5]. Since $E_{cu}^{*}\cap E_{cs}^{*}=\emptyset$, this means that the distributional restriction $\theta_{x}$ where for any $g\in C^{\infty}(\mathbb{T}^{3})$, $\theta_{x}(g)=(\theta,g[\mathcal{W}^{cs}(x)])_{\mathcal{H}^{s}\times\mathcal{H}^{-s}}$ is well defined [29, Corollary 4.2.2]. We first show that $(\theta_{x})_{x\in\mathbb{T}^{3}}$ defines a system of $cs$-measures.
@@
Since $\mathrm{WF}(g[\mathcal{W}^{cs}(x)])\subset E_{cs}^{*}\subset E_{s}^{*}$, one has $g[\mathcal{W}^{cs}(x)]\in\mathcal{H}^{-s}$ for some $s>0$, see [20, Lemma 1.11]. Lemma 4.7 then gives, for some $c>0$,
@@
where we used (1.6) for the resonance $\lambda_{u}$. We now show the invariance by unstable holonomy. The argument is given in [9, Section 4.1] but we include the proof here for completeness. Note that since $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$, $f^{-1}$ is a partially hyperbolic diffeomorphism with dominated splitting $T(\mathbb{T}^{3})=E_{u}\oplus E_{cs}$ with $E_{u}$ being contracting for $f^{-1}$. We can thus use the absolute continuity of $E_{u}$, see [9, Theorem 4.3]. Namely, there are constants $K,\beta>0$ such that for any $\varepsilon>0$, for any rectangle $R$, for any $x_{1},x_{2}\in R$ and for any two $\varepsilon$-equivalent Borel sets $A_{i}\subset\mathcal{W}^{cs}(x_{i})$, $i=1,2$ (see §2.1.3), the measures $(\mathrm{Hol}^{u}_{x_{2},x_{1}})_{*}\mathrm{vol}_{A_{1}}$ and $\mathrm{vol}_{A_{2}}$ are equivalent and

### v3

@@ 3. Construction of a Margulis system of measures
which shows that $(\tilde{\mu}^{u}_{x})_{x\in\mathbb{T}^{3}}$ has full support in each leaf and is finite on compact sets. Note that it is not clear at this point that $(\tilde{\mu}^{u}_{x})_{x\in\mathbb{T}^{3}}$ is measurable.
A definition of the system of measures using dimension theory. We now define the Margulis system $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ using the system of measures $(\tilde{\mu}^{u}_{x})_{x\in\mathbb{T}^{3}}$ constructed above and using ideas from dimension theory, following Climenhaga, Pesin et Zelerowicz [undefk]. This more explicit construction will allow us to show that the system is measurable. We recall the definition of a $u$-Bowen ball:
(3.15) $\forall x\in\mathbb{T}^{3},\ \forall n\geq 0,\ \forall\delta>0,\quad\mathcal{W}^{u}_{n}(x,\delta):=\{y\in\mathbb{T}^{3}\mid\max_{k=0}^{n}d^{u}(f^{k}x,f^{k}y)<\delta\}.$
Since we work with an adapted metric, notice that $\mathcal{W}^{u}_{n}(x,\delta)=f^{-n}\mathcal{W}^{u}(f^{n}(y),\delta)$. For a Borel set $Z\subset\mathcal{W}^{u}(x)$, the set of open covers by $u$-Bowen balls is
(3.16) $\forall N\geq 0,\quad\mathcal{C}_{N}(Z):=\{\mathcal{C}\subset\mathbb{T}^{3}\times\llbracket N,+\infty\llbracket,\ Z\subset\cup_{(y,n)\in\mathcal{C}}\mathcal{W}^{u}_{n}(y,\delta)\}.$
The system of measures is defined as follows:
(3.17) $\mu^{u,N}_{x}(Z):=\inf_{\mathcal{C}\in\mathcal{C}_{N}(Z)}\sum_{(y,n)\in\mathcal{C}}\lambda_{u}^{-n},\quad\mu^{u}_{x}(Z):=\lim_{N\to+\infty}\mu^{u,N}_{x}(Z).$
For any $x\in\mathbb{T}^{3}$, $\mu^{u}_{x}$ defines a Borelian measure on $\mathcal{W}^{u}(x)$. It is clear from (3.17) that $\mu^{u}_{x}(Z)\in[0,+\infty]$ for any Borel subset $Z\subset\mathcal{W}^{u}(x)$. We show that $\mu^{u}_{x}$ is an metric outer measure. Recall that a measure $m$ on a metric space $(X,d)$ is a metric outer measure if for any $A,B\subset X$ such that $d(A,B):=\inf\{d(a,b)\mid a\in A,b\in B\}>0$, one has $m(A\cup B)=m(A)+m(B)$. Every metric outer measure has the property that all Borel sets are measurable, see for instance [undefah, Proposition 12.41].
Suppose that $Z_{1},Z_{2}\subset\mathcal{W}^{u}(x)$ are such that $d^{u}(Z_{1},Z_{2})>0$. Since the $u$-Bowen balls shrink exponentially fast when $N\to+\infty$, for $N$ large enough, $\mathcal{C}_{N}(Z_{1}\cup Z_{2})$ splits as a part that covers $Z_{1}$ and a part that covers $Z_{2}$. In particular, one has
$\forall N\gg 1,\ \mu^{u,N}_{x}(Z_{1}\cup Z_{2})=\mu^{u,N}_{x}(Z_{1})+\mu^{u,N}_{x}(Z_{2})\ \Rightarrow\ \mu^{u}_{x}(Z_{1}\cup Z_{2})=\mu^{u}_{x}(Z_{1})+\mu^{u}_{x}(Z_{2}).$
This shows that $\mu^{u}_{x}$ is a Borel measure on $\mathcal{W}^{u}(x)$.
The measure $\mu^{u}_{x}$ is finite on compact sets. Let $K\subset\mathcal{W}^{u}(x)$ be a compact subset. Let $N\geq 0$ and consider $\mathcal{C}\in\mathcal{C}_{N}(K)$ such that no proper subcover of $\mathcal{C}$ is in $\mathcal{C}_{N}(K)$. One has
$K\subset\bigcup_{(y,n)\in\mathcal{C}}\mathcal{W}^{u}_{n}(y,\delta)=\bigcup_{(y,n)\in{\mathcal{C}}}f^{-n}\big(\mathcal{W}^{u}(f^{n}y,\delta)\big).$
Since $\mathcal{W}^{u}(x)$ is one dimensional and since no proper subcover of $\mathcal{C}$ is in $\mathcal{C}_{N}(K)$, for any $z\in K$, there are at most two elements $(y,n)\in{\mathcal{C}}$ such that $z\in\mathcal{W}^{u}_{n}(y,\delta)$. Hence,
$\sum_{(y,n)\in\mathcal{C}}\tilde{\mu}^{u}_{x}\big(f^{-n}\big(W^{u}(f^{n}y,\delta)\big)\big)\leq 2\tilde{\mu}^{u}_{x}(K).$
Using (3.13) and (3.14), we obtain
$\sum_{(y,n)\in\mathcal{C}}\lambda_{u}^{-n}\leq 2C_{1}(\delta)^{-1}\tilde{\mu}^{u}_{x}(K).$
This shows that $\mu^{u}_{x}(K)\leq 2C_{1}(\delta)^{-1}\tilde{\mu}^{u}_{x}(K)$ is finite.
The measure $\mu^{u}_{x}$ has full support in $\mathcal{W}^{u}(x,\delta)$. Let $x\in\mathbb{T}^{3}$ and $\varepsilon>0$. We show that $\mu^{u}_{x}(\mathcal{W}^{u}(x,\varepsilon))>0$. Let $\mathcal{C}\in\mathcal{C}_{N}(\mathcal{W}^{u}(x,\varepsilon))$. Since $\mathcal{W}^{u}(x,\varepsilon)$ is relatively compact, we can suppose without loss of generality that $\mathcal{C}$ is finite. We have
$\mathcal{W}^{u}(x,\varepsilon)\subset\bigcup_{(y,n)\in\mathcal{C}}\mathcal{W}^{u}_{n}(y,\delta)=\bigcup_{y\in\mathcal{C}}f^{-n}\big(\mathcal{W}^{u}(f^{n}y,\delta)\big).$
This implies that
$\tilde{\mu}^{u}_{x}(\mathcal{W}^{u}(x,\varepsilon))\leq\sum_{(y,n)\in\mathcal{C}}\tilde{\mu}^{u}_{x}\big(f^{-n}\big(\mathcal{W}^{u}(f^{n}y,\delta)\big)\big).$
Using (3.13) and (3.14), we obtain
$C_{1}(\varepsilon)C_{2}(\delta)^{-1}\leq\sum_{(y,n)\in\mathcal{C}}\lambda_{u}^{-n}.$
Passing to the infimum gives $\mu^{u}_{x}(\mathcal{W}^{u}(x,\varepsilon))\geq C_{1}(\varepsilon)C_{2}(\delta)^{-1}>0$.
The Margulis system has the scaling property (3.8). Let $Z\subset\mathcal{W}^{u}(x)$ be a relatively compact Borel set. We need to show that $\mu^{u}_{f(x)}(f(Z))=\lambda_{u}\mu^{u}_{x}(Z)$. Let $N\geq 0$ be an integer and $\mathcal{C}\in C_{N}(Z)$. Then if we define $\mathcal{C}^{\prime}:=\{(f(x),n-1)\mid(x,n)\in\mathcal{C}_{N}(Z)\}$, we see that $\mathcal{C}^{\prime}\in\mathcal{C}_{N-1}(f(Z))$. Moreover, any $\mathcal{C}^{\prime}\in\mathcal{C}_{N-1}(f(Z))$ can be obtained this way and
$\sum_{(y,m)\in\mathcal{C}^{\prime}}\lambda_{u}^{-m}=\sum_{(z,n)\in\mathcal{C}}\lambda_{u}^{-(n-1)}=\lambda_{u}\sum_{(z,n)\in\mathcal{C}}\lambda_{u}^{-n}.$
Taking the infimum yields $\mu^{u}_{f(x)}(f(Z))=\lambda_{u}\mu^{u}_{x}(Z)$.
The Margulis system $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ is measurable. Since the limit of a sequence of measurable functions is measurable, (3.17) shows that it suffices to prove that $(\mu^{u,N}_{x})_{x\in\mathbb{T}^{3}}$ is measurable. Let $M\in\mathbb{N}$ with $M>N$. Define
$\mathcal{C}_{N,M}(Z):=\{\mathcal{C}\subset\mathbb{T}^{3}\times\llbracket N,M\rrbracket,\ Z\subset\cup_{(y,n)\in\mathcal{C}}\mathcal{W}^{u}_{n}(y,\delta)\}$
as well as
(3.18) $\mu^{u,N,M}_{x}(Z):=\inf_{\mathcal{C}\in\mathcal{C}_{N,M}(Z)}\sum_{(x,n)\in\mathcal{C}}\lambda_{u}^{-n}.$
Since $\mu^{u,N}_{x}:=\lim_{M\to+\infty}\mu^{u,N,M}_{x}$, we only need to show that $(\mu^{u,N,M}_{x})_{x\in\mathbb{T}^{3}}$ is measurable. To that end, first consider a compact set $K\subset\mathbb{T}^{3}$. We prove that $x\mapsto\mu^{u,N,M}_{x}(K_{x})$, where $K_{x}:=K\cap\mathcal{W}^{u}_{\mathrm{loc}}(x)$, is upper semi-continuous, thus measurable. By compactness of $K$, $x\mapsto K_{x}$ is upper semicontinuous. For our purpose, this means that for any $x\in\mathbb{T}^{3}$ and any $\varepsilon>0$, there is a $\eta>0$ such that if $d(x,y)<\eta$, then
$K_{y}\subset\mathrm{Hol}^{cs}_{x,y}(U_{\varepsilon}(K_{x}))$
where $U_{\varepsilon}(K_{x}):=\cup_{z\in K_{x}}\mathcal{W}^{u}(z,\varepsilon)$ is the $\varepsilon$-neighborhood of $K_{x}$. Let $\mathcal{C}\in\mathcal{C}_{N,M}(K_{x})$ with no proper subcover in $\mathcal{C}_{N,M}(K_{x})$. This means that
$K_{x}\subset\bigcup_{(y,n)\in\mathcal{C}}f^{-n}\big(\mathcal{W}^{u}(f^{n}y,\delta)\big).$
As before, this implies
$\sum_{(y,n)\in\mathcal{C}}\tilde{\mu}^{u}_{x}\big(f^{-M}\big(\mathcal{W}^{u}(f^{n}y,\delta)\big)\big)\leq\sum_{(y,n)\in\mathcal{C}}\tilde{\mu}^{u}_{x}\big(f^{-n}\big(\mathcal{W}^{u}(f^{n}y,\delta)\big)\big)\leq 2\tilde{\mu}^{u}_{x}(K).$
Using (3.13) and (3.14), we obtain $|\mathcal{C}|\leq 2C_{1}(\delta)^{-1}\tilde{\mu}^{u}_{x}(K_{x})\lambda_{u}^{M}.$ In particular, the infimum in (3.18) can be taken over partitions of uniformly bounded cardinality. As a consequence, the function in (3.18) takes its values in a finite set and thus attains its minimum. We now suppose that $\mathcal{C}\in\mathcal{C}_{N,M}(K_{x})$ is such that
$\mu^{u,N,M}_{x}(K_{x})=\sum_{(y,n)\in\mathcal{C}}\lambda_{u}^{-n}.$
Since $K_{x}$ is closed, there exists $\varepsilon>0$ and $\delta^{\prime}<\delta$ such that
$U_{\varepsilon}(K_{x})\subset\bigcup_{(y,n)\in\mathcal{C}}f^{-n}\big(W^{u}(f^{n}y,\delta^{\prime})\big).$
Let $y\in\mathbb{T}^{3}$ be $\eta$-close to $x$ where $\eta$ is defined by the upper semi-continuity of $K$. Up to taking $\eta$ even smaller and since all the foliations are continuous, we have
$\mathrm{Hol}^{cs}_{f^{n}x,f^{n}y}\big(W^{u}(f^{n}z,\delta^{\prime})\big)\subset\mathcal{W}^{u}(\mathrm{Hol}^{cs}_{f^{n}x,f^{n}y}\circ f^{n}z,\delta)=\mathcal{W}^{u}(f^{n}\circ\mathrm{Hol}^{cs}_{x,y}(z),\delta).$
In other words, we have shown that $\mathcal{C}^{\prime}:=\{(\mathrm{Hol}^{cs}_{x,y}(z),n)\mid(z,n)\in\mathcal{C}\}$ is an element of $\mathcal{C}_{N,M}(K_{y})$. Moreover, one has
$\mu^{u,N,M}_{x}(K_{x})=\sum_{(z,n)\in\mathcal{C}}\lambda_{u}^{-n}=\sum_{(w,n)\in\mathcal{C}^{\prime}}\lambda_{u}^{-n}.$
Passing to the infimum, we thus have $\mu^{u,N,M}_{y}(K_{y})\leq\mu_{x}^{u,N,M}(K_{x})$ which finishes the proof for compact sets $K$.
Now, let $\mathcal{K}$ be the collection of all subsets $Z\subset\mathbb{T}^{3}$ such that $x\mapsto\mu^{u,N,M}_{x}(Z_{x})$ is measurable. Note that this is a Dynkin system, i.e.,
  • If $Z_{1},Z_{2}\in\mathcal{K}$ and $Z_{1}\subset Z_{2}$, then $Z_{2}\setminus Z_{1}\in\mathcal{K}$.
  • If $Z_{1}\subset Z_{2}\subset\ldots\subset Z_{n}\subset\ldots\in\mathcal{K}$, then $\cup_{n\in\mathbb{N}}Z_{n}\in\mathcal{K}$.
  • $\mathbb{T}^{3}\in\mathcal{K}$.
Since $\mathcal{K}$ contains the closed subsets, Dynkin $\pi$-$\lambda$ theorem implies that it must contain the $\sigma$-algebra generated by the closed subsets. In other words, $\mathcal{K}$ contains all Borel sets which concludes the proof. ∎
@@
Having constructed a measurable system of $u$-measures, we now characterize $u$-MMEs by their conditionals along the strong unstable foliation. We recall the following result due to Buzzi, Fisher and Tazhibi [undefh], see also [undefam, Proposition 5.3].
@@ 4.1. Ruelle resonances
In this subsection, we recall some facts about the functional approach for Anosov diffeomorphisms. For a detailed introduction to the subject, we refer to [undefa]. For $0\leq k\leq 3$, define the vector bundle of smooth $k$-forms $\Omega_{k}=C^{\infty}(\mathbb{T}^{3};\Lambda^{k}T^{*}(\mathbb{T}^{3})).$ The diffeomorphism $f$ acts on $\Omega_{k}$ by pushforward:
@@
For an Anosov diffeomorphism, one can associate to $f_{*,k}$ a discrete spectrum $\mathrm{Res}_{k}(f)$, the Pollicott-Ruelle resonances, by making it act on specially designed anisotropic spaces, see for instance [undefd, undefb, undefq, undefo].
@@
where $D^{\prime}_{k}(\mathbb{T}^{3})$ denotes the space of distributional currents of degree $k$, i.e., the dual space of $\Omega_{3-k}$, see [undefab, Chapter 10.1]. We will write $(\cdot,\cdot)_{\Omega_{k} \times\mathcal{D}^{\prime}_{3-k}}$ for the pairing between $\Omega_{k}$ and $\mathcal{D}^{\prime}_{3-k}$. More precisely, by the work of Faure, Roy and Sjöstrand [undefo], there exists a family of Hilbert spaces $(\mathcal{H}^{s}(\mathbb{T}^{3};\Omega_{k}))_{s>0}$ such that:
  • the space $\Omega_{k}$ is densely included in $\mathcal{H}^{s}(\mathbb{T}^{3};\Omega_{k})$ (see [undefab, Lemma 9.1.13]) and one has $H^{s}\subset\mathcal{H}^{s}\subset H^{-s}$, where $H^{s}$ is the usual $L^{2}$-Sobolev space of order $s$ (see [undefab, Lemma 9.1.14]);
@@
Applying the result to $f^{-1}$, we obtain a family of Hilbert spaces $(\mathcal{H}^{-s}(\mathbb{T}^{3},\Omega_{k}))_{s>0}$, dual to $(\mathcal{H}^{s}(\mathbb{T}^{3},\Omega_{k}))_{s>0}$ on which the resolvent of the pullback of $f$ has a meromorphic extension, see [undefo]. We will write $(\cdot,\cdot)_{\mathcal{H}^{s} \times\mathcal{H}^{-s}}$ for the pairing between $\mathcal{H}^{s}$ and $\mathcal{H}^{-s}$.
@@
where $R_{k}^{H}(\lambda)$ is the holomorphic part near $\lambda_{0}$. The generalized resonant states are the elements in the range of the spectral projector.
@@
where $\mathrm{WF}(u)$ denotes the wavefront set of a current $u$, see [undefv, Chapter VIII]. The resonant states are defined as
@@ 4.2. Zeta function
The zeta function $\zeta_{f}$ has a meromorphic extension (actually rational extension in our case) to the complex plane. This is a special case of the works of Liverani-Tsujii [undefac] and Baladi-Tsujii [undefc]. In fact, for an Anosov diffeomorphism on a torus, one can compute explicitly $\zeta_{f}$ in terms of its action on the homology $A$.
@@ 4.3. Dynamical determinants
We will need the following result, see for instance [undefa, Theorem 6.2]. For any $k\in\llbracket 0,3\rrbracket$, the dynamical determinant $\mathcal{D}_{f,k}$ admits a holomorphic extension to $\mathbb{C}$. Moreover,
@@
We recognize the Ruelle zeta function with potential $J^{c}_{f}$. Since $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ is conjugate to a linear automorphism of $\mathbb{T}^{3}$, it is topologically mixing. By [undefai, Theorem 1], there is a $R=R(J^{c}_{f})>e^{-P(J^{c}_{f})}$ such that $\zeta_{f,J^{c}_{f}}$ has no zero in $\overline{B}(0,R)$ except for a simple zero at $e^{-P(J^{c}_{f})}$. This shows that $e^{P(J^{c}_{f})}$ is a resonance of algebraic multiplicity equal to $1$ and that $e^{P(J^{c}_{f})}$ is the only resonance on the critical circle. ∎
@@
see [undefa, Equation $(6.38)$]. Using (4.9), we also show the following.
@@
We show that the spectral projector at the first resonance is equal to $\mu_{J^{c}_{f}}$.
@@
Let $g\in C^{\infty}(\mathbb{T}^{3},\mathbb{R}_{+})$ be a smooth and positive function. The Guillemin trace formula (see for instance [undefc, Proposition 6.3]) gives for any $m\geq 0$,
@@
This means that (4.17) is Bowen’s formula for the equilibrium state (see for instance $\mu_{J^{c}_{f}}$ [undeft, Theorem 20.3.7]) and this concludes the proof. ∎
@@
Let $\omega\in\mathcal{H}^{s}$, then there is $c_{\omega}\in\mathbb{R}$ such that $\Pi_{0}(\omega)=c_{\omega}\theta$. We obtain (4.18) by applying the previous convergence to $\omega$. The proof of (4.19) is analogous.
Remark 2.
The resonant ((resp. co-resonant)) state $\theta$ ((resp. $\nu)$ is closed. Indeed, since $d$ commutes with the pushforward (resp. pullback) and preserves the spaces of currents with wavefront set in $E_{s}^{*}$ ((resp. $E_{u}^{*})$, one has $d\theta\in\mathrm{Res}_{3,\lambda_{u}}(f)$ ((resp. $d\nu\in\mathrm{Res}^{*}_{2,\lambda_{u}}(f))$. A similar argument as in the proof of Lemma 4.5 shows that $\mathrm{Res}_{3}(f)\subset\{z\in\mathbb{C},\ |z|\leq 1\}$ which then implies that $d\theta=0$. For the co-resonant state $\nu$, we can use (4.15) with (4.11) to conclude that $\mathrm{dim}(\mathrm{Res}^{*}_{1,\infty,\lambda_{u}}(f))-\mathrm{dim}(\mathrm{Res}^{*}_{2,\infty,\lambda_{u}}(f))=1$. Since Lemma 4.4 shows that $\mathrm{dim}(\mathrm{Res}^{*}_{1,\infty,\lambda_{u}}(f))=1$, we deduce that $\mathrm{Res}^{*}_{2,\infty,\lambda_{u}}(f))=\{0\}$ and thus $d\nu=0$. This should be compared with the work of Ruelle and Sullivan on the system of Margulis measures for an Anosov diffeomorphism, see [undefaj, Corollary p.326].
@@
By [undeft, Theorem 20.3.9], $\mu_{J^{c}_{f}}$ is a measure of maximal entropy if and only if $J^{c}_{f}$ is cohomologous to a constant. By the Livsic theorem [undeft, Theorem 19.2.1], this holds if and only if for any periodic point $p$ of period $n(p)$, one has
@@
The fact that the constant has to be equal to $\lambda_{c}$ follows from [undefp, Corollary 2.6]. This holds if and only if $E_{u}$ and $E_{s}$ are jointly integrable by [undefp, Theorem 5.1]. ∎
@@
We note that by Corollary 1.2 and Corollary 4.2, the system of $u$-measures $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ coincides (up to a constant rescaling) with the conditionals $(\mu^{u}_{J^{c}_{f},x})_{x\in\mathbb{T}^{3}}$ of $\mu_{J^{c}_{f}}$ along $\mathcal{W}^{u}$. We first start by expressing the $cs$-conditionals $(\mu^{cs}_{J^{c}_{f},x})_{x\in\mathbb{T}^{3}}$ using the resonant state $\theta$. By (1.6), one has $\mathrm{WF}(\theta)\subset E_{cu}^{*}$ and for any $x\in\mathbb{T}^{3}$, $\mathrm{WF}([\mathcal{W}^{cs}(x)])\subset E_{cs}^{*}$, where $\mathcal{W}^{cs}(x)]$ denotes the integration current on $\mathcal{W}^{cs}(x)$, see for instance [undefab, Example 4.1.5]. Since $E_{cu}^{*}\cap E_{cs}^{*}=\emptyset$, this means that the distributional restriction $\theta_{x}$ where for any $g\in C^{\infty}(\mathbb{T}^{3})$, $\theta_{x}(g)=(\theta,g[\mathcal{W}^{cs}(x)])_{\mathcal{H}^{s}\times\mathcal{H}^{-s}}$ is well defined [undefab, Corollary 4.2.2]. We first show that $(\theta_{x})_{x\in\mathbb{T}^{3}}$ defines a system of $cs$-measures.
@@
Since $\mathrm{WF}(g[\mathcal{W}^{cs}(x)])\subset E_{cs}^{*}\subset E_{s}^{*}$, one has $g[\mathcal{W}^{cs}(x)]\in\mathcal{H}^{-s}$ for some $s>0$, see [undefs, Lemma 1.11]. Lemma 4.7 then gives, for some $c>0$,
@@
where we used (1.6) for the resonance $\lambda_{u}$. We now show the invariance by unstable holonomy. The argument is given in [undefh, Section 4.1] but we include the proof here for completeness. Note that since $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$, $f^{-1}$ is a partially hyperbolic diffeomorphism with dominated splitting $T(\mathbb{T}^{3})=E_{u}\oplus E_{cs}$ with $E_{u}$ being contracting for $f^{-1}$. We can thus use the absolute continuity of $E_{u}$, see [undefh, Theorem 4.3]. Namely, there are constants $K,\beta>0$ such that for any $\varepsilon>0$, for any rectangle $R$, for any $x_{1},x_{2}\in R$ and for any two $\varepsilon$-equivalent Borel sets $A_{i}\subset\mathcal{W}^{cs}(x_{i})$, $i=1,2$ (see §2.1.3), the measures $(\mathrm{Hol}^{u}_{x_{2},x_{1}})_{*}\mathrm{vol}_{A_{1}}$ and $\mathrm{vol}_{A_{2}}$ are equivalent and